# references.bib

### v1

@@ 2.2 Two-step Floquet dynamics
Here we show how to fix the values of the inhomogeneity parameters $\mathbf{w}$ of the transfer matrix for both the periodic and open boundary conditions, in order to define an integrable discrete time dynamics. The original idea comes from the work by Baxter [baxter2016exactly] where the value of the inhomogeneities are fixed to define a discrete time parallel update on the periodic lattice. Our construction initially follows [vanicat2018integrable, vanicat2018integrable2]. In this work, our main focus is the case of open boundary conditions. For the reasons that will be clear in the following, we construct a quantum circuit where each gate is composed by the tensor product of two $R$-matrices of a spin 1/2 chain. Due to the tensor product structure of local Hilbert space, it is easy to see that a (trivial) solution of the boundary reflection algebra can be found by taking the tensor product of two solutions of the reflection algebra for the single $R$-matrix, see Appendix A. We explore the non-trivial cases where the solutions of the boundary reflection algebra in the enlarged Hilbert space is richer.
@@ 2.2.1 Periodic boundary condition
$w_{2n}=\kappa,\qquad w_{2n+1}=-\kappa.$ (2.21)
The transfer matrix (2.4) is
$t(u)=\text{tr}_{0}\Big{(}R_{0L}(u-\kappa)R_{0,L-1}(u+\kappa)\dots R_{02}(u-\kappa)R_{01}(u+\kappa)\Big{)}.$ (2.22)
@@
$U_{i,j}=\check{R}_{i,j}(2\kappa)=P_{i,j}R_{i,j}(2\kappa).$ (2.23)
@@
$t(\kappa)=$ $P_{1,L}P_{1,L-1}\dots P_{1,3}P_{1,2}U_{2,3}\dots U_{L-4,L-3}U_{L-2,L-1}U_{L,1}$ (2.24)
@@
$M=t(-\kappa)^{-1}t(\kappa)=\mathbb{U}^{o}\mathbb{U}^{e},$ (2.26)
@@
$\mathbb{U}^{o}=\prod_{k=1}^{L/2}U_{2k-1,2k},$ $\mathbb{U}^{e}=\prod_{k=1}^{L/2}U_{2k,2k+1}.$ (2.27)
The operator $M$ describes one period of Floquet dynamics composed of two steps: $\mathbb{U}^{e}$ and $\mathbb{U}^{o}$. $\mathbb{U}^{e}$ updates every pair of consecutive sites with the left site at even positions, while $\mathbb{U}^{o}$ with the left site at odd positions [^7] We use the convention to enumerate the sites from right to left., see Fig. 3.
@@
This construction builds an integrable quantum circuit. In fact, since $[t(u,\mathbf{w}),t(v,\mathbf{w})]=0,$ fixing the inhomogeneities to (2.21), it is easy to see that
$[M,t(z)]=0,$ (2.28)
with $t(z)$ given in (2.22).
@@ 2.2.2 Open boundary conditions
$w_{1}=w_{3}=w_{5}=\dots=w_{L}=\kappa,$ $w_{2}=w_{4}=\dots=w_{L-1}=-\kappa.$ (2.29)
The transfer matrix (2.15) is now
$t(u)=\text{tr}_{0}\Big{(}{K}_{0}^{L}(u)R_{0L}(u\!-\!\kappa)\dots R_{02}(u\!+\!\kappa)R_{01}(u\!-\!\kappa)K_{0}^{R}(u)R_{10}(u\!+\!\kappa)R_{20}(u\!-\!\kappa)\dots R_{L0}(u\!+\!\kappa)\Big{)}.$ (2.30)
@@
$t(\kappa)=K_{1}^{R}(\kappa)U_{23}U_{45}\dots U_{L-1,L}U_{12}U_{34}\dots U_{L-2,L-1}\bar{K}_{L}^{L}(\kappa),$ (2.31)
@@
$\bar{K}_{L}^{L}(\kappa)=\text{tr}_{0}\big{(}K_{0}^{L}(\kappa)R_{0L}(2\kappa)P_{0L}\big{)}.$ (2.32)
@@
$M=t(\kappa)=\mathbb{U}^{e}\mathbb{U}^{o},$ (2.33)
@@
$\mathbb{U}^{o}=\bar{K}_{L}^{L}(\kappa)\left(\prod_{k=1}^{\frac{L-1}{2}}U_{2k-1,2k}\right)\,,$ $\mathbb{U}^{e}=\left(\prod_{k=1}^{\frac{L-1}{2}}U_{2k,2k+1}\right)K^{R}_{1}(\kappa).$ (2.34)
@@ 3 Our construction
In this section, we first discuss in 3.1, the construction of the quantum circuit we are going to study and we clarify the main question we would like to address. In 3.2, we discuss the methods.
@@ 3.1 Building the quantum circuit
$U_{ik,jl}=\check{R}_{ik,jl}(u)=\check{r}_{ij}(u)\check{r}_{kl}(-u).$ (3.1)
We represent this "double gate" as in Fig. 5, [bertini2020operator].
@@
For open boundary condition, a natural generalization of the right and left reflection algebra is to consider $k^{R/L}(u)$ and $\tilde{k}^{R/L}(u)$ as two solutions of the boundary Yang-Baxter equation (2.9) and (2.10) for $r(u)$ and construct the solution of the boundary Yang-Baxter equation for $R(u)$ as $K^{R/L}_{ij}(u)=k^{R/L}_{i}(u)\tilde{k}^{R/L}_{j}(-u)$, see Appendix A for a proof of this statement. The two-steps Floquet circuit for this construction is Figure 7.
@@ 3.2 Key steps
  • Solve the reflection equation (2.9) for $K^{R}(u)$;
  • Solve the reflection equation [^8] This solution can be found by applying to the $K^{R}$ solution the automorphism (2.20). If the $r(u)$ of the model analyzed obeys the crossing unitarity property, this automorphism reduce to the three (equivalent) given in section 2.1.2. (2.10) for $K^{L}(u)$;
@@
To solve the equation (2.9), we consider the most general form for the $K^{R}$-matrix, as a $4\times 4$ matrix with entries being functions dependent on a spectral parameter $u$. We label the entries of the $K^{R}$ matrix $\mathcal{K}_{i,j}(u)$, with $i,j$ identifying the row or the column where this element is placed. The expression of the $R$-matrix depends on the model we are analyzing. We normalize [^9] Two $K$ matrices that differ only by a normalization factor solve the same boundary YBE. We choose $\mathcal{K}_{2,2}=1$ since we require that $K^{R}(0)=\mathbb{I}$. Choosing to normalize an off-diagonal element to one won’t be compatible with the regularity condition. the element $\mathcal{K}_{2,2}$ to one. A possible strategy to solve the boundary Yang-Baxter equation is to use a differential approach called Abel’s method. For a review of the method, see [Vieira:2017vnw]. This method consists of taking the formal derivatives of the equation (2.9) with respect to the spectral parameters $u$, then evaluate it at zero and use the regularity condition
$K^{R}(0)=\mathbb{I},$ $\mathcal{K}_{i,j}(0)=\delta_{i,j}.$ (3.2)
@@
We take the formal derivative of the boundary Yang-Baxter equation (2.9),
@@
In this way, we obtain a system of equations depending on the variables $\mathcal{K}_{i,j}(v)$ and $\kappa_{i,j}$. The advantage to solve this system, compared to the initial one given by (2.9), is that now the equations are linear in the variables. To find all the solutions, we treat $\mathcal{K}_{i,j}$ as independent from $\kappa_{i,j}$. We solved some of the equations [^10] The choice on how many equations to solve in the first place depends on the model considered. So far, we did not find a systematic way to make this choice. for some of the variables and plugging back into the remaining equations we solved for the others. Last, we fixed the remaining unknown functions by imposing the compatibility conditions that ${\mathcal{K}_{i,j}}^{\prime}(0)=\kappa_{i,j}$. After having found all the solutions, we discard the ones leading to incompatible conditions [^11] An example of incompatible conditions may be that the element $\mathcal{K}_{2,2}=3$ and $\kappa_{2,2}(0)=5$.. We also discard the solutions that are not compatible with the regularity conditions.
This method is general and only requires that the functions $\mathcal{K}_{i,j}(u)$ are differentiable. This condition, for our case, is not restrictive at all since, as explained in section 5, the $K^{R}$ is one of the building block of the transfer matrix and hence, its derivatives correspond to the conserved charge characterizing the integrable model. Furthermore, (2.9) contains both the $u$ and $v$, spectral parameter dependence, taking the derivative w.r.t. $u$ and later sending it to zero, won’t imply that we are finding solutions only at order linear in $u$.
This method was vastly used in the literature, in particular Lima-Santos and collaborators, classified the $K$-matrices associated with non-exceptional Lie algebras and superalgebras, see for example [malara2006and, vieira2017reflection].
@@ 4.1.1 Heisenberg (XXX) spin chain
We solve the boundary Yang-Baxter equation (2.9) for $K^{R}$(u) and
$R_{ij,kl}(u)=\big{(}{p}_{ik}+u\mathbb{I}\big{)}\big{(}{p}_{jl}-u\mathbb{I}\big{)}.$ (4.2)
We solved the equation (2.9), as explained in the previous section, by using the Mathematica software and we obtained all the solutions for the right reflection matrix $K^{R}(u)$. After detailed investigation, we found that all the solutions of the boundary YBE for two copies of the Heisenberg spin chain take the factorized form
${K}_{ij}^{R}(u)=k^{R}_{i}(u)\tilde{k}^{R}_{j}(-u).$ (4.3)
Both $k_{i}(u)$ and $\tilde{k}_{j}(u)$ are solutions of the boundary YBE (2.9) for $r(u)$.
The Heisenberg spin chain is widely studied, diagonal solutions were classified in [Sklyanin:1988yz] and general solutions in [de1994boundary]. For completeness, we write here [^13] This solution is related to the one of [de1994boundary] by a reparametrization. the most general solution of the reflection algebra for the XXX spin chain with $k^{R}$ regular,
$k^{R}(u)=\left(\begin{array}[]{cc}\frac{\kappa_{1}u+1}{1-\kappa_{1}u}&\frac{\kappa_{2}u}{1-\kappa_{1}u}\\ \frac{\kappa_{4}u}{1-\kappa_{1}u}&1\\ \end{array}\right),$ (4.6)
where $\kappa_{1}$, $\kappa_{2}$ and $\kappa_{4}$ are free constants. To construct the most general $K^{R}(u)$ from (4.3) , one can assume that $\tilde{k}_{j}(u)$ has the same form as (4.6) and relabel the constants $\kappa_{i}$ as $\tilde{\kappa}_{i}$ since now they can in principle take different values. This is described in more details in the Appendix A.
To obtain the solution of the left boundary $K^{L}$ one can use the automorphism (2.20). This guarantees that if the solution for $K^{R}$ can be written in the factorized form, then the one for $K^{L}$ also factorizes. For this model, we used the more general automorphism (2.20), because the $R$-matrix $R(u)$ (4.2) does not have the crossing unitarity property. For the Heisenberg spin chain, the crossing parameter for $r(u)$ is $\eta=1$, so the crossing unitarity property is broken for $R(u)=r(u)\otimes r(-u)$. However, since we know from the more general automorphism that the solutions for $K^{L}$ should also factorize, we can separately obtain the solution for $k^{L}(u)$ and $\tilde{k}^{L}(u)$. In this way, we obtain
$K^{L}(u)=(k^{R})^{t}(-u-1)\otimes(\tilde{k}^{R})^{t}(u-1).$ (4.7)
@@ 4.1.2 XXZ spin chain
$r(u)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\sin(u)\csc(u-i\gamma)&-i\sinh(\gamma)\csc(u-i\gamma)&0\\ 0&-i\sinh(\gamma)\csc(u-i\gamma)&\sin(u)\csc(u-i\gamma)&0\\ 0&0&0&1\\ \end{array}\right),$ (4.12)
@@
$k(u)=\left(\begin{array}[]{cc}\frac{\kappa_{1,1}\sin u+2\cos u}{2\cos u-\kappa_{1,1}\sin u}&\frac{2\kappa_{1,2}}{2\csc u-\kappa_{1,1}\sec u}\\ \frac{2\kappa_{2,1}}{2\csc u-\kappa_{1,1}\sec u}&1\\ \end{array}\right)$ (4.16)
and $\tilde{k}(u)$ is the same by substituting $\kappa$ with $\tilde{\kappa}$. This solution reproduces (4.6), in the same spirit as one gets the rational $R$-matrix of the Heisenberg spin chain from the trigonometric solution for the XXZ spin chain, [analecture]. As observed in [de1994boundary], the solution of the right boundary Yang Baxter equation is independent from the parameter $\gamma$. However, from the automorphisms (2.20) it is clear that $K^{R}$ is connected to $K^{L}$ via the $R$-matrix and hence the anisotropy $\Delta$ (or $\gamma$) of the spin chain enters. We can repeat the same argument of the Heisenberg spin chain, and we obtain
$K^{L}(u)=(k^{R})^{t}(-u+i\gamma)\otimes(\tilde{k}^{R})^{t}(u+i\gamma).$ (4.17)

### v2

@@ 2.2 Two-step Floquet dynamics
Here we show how to fix the values of the inhomogeneity parameters $\mathbf{w}$ of the transfer matrix for both the periodic and open boundary conditions, in order to define an integrable discrete time dynamics. The original idea comes from the work by Baxter [7] where the value of the inhomogeneities are fixed to define a discrete time parallel update on the periodic lattice. Our construction initially follows [8, 9]. In this work, our main focus is the case of open boundary conditions. For the reasons that will be clear in the following, we construct a quantum circuit where each gate is composed by the tensor product of two $R$-matrices of a spin 1/2 chain. Due to the tensor product structure of local Hilbert space, it is easy to see that a (trivial) solution of the boundary reflection algebra can be found by taking the tensor product of two solutions of the reflection algebra for the single $R$-matrix, see Appendix A. We explore the non-trivial cases where the solutions of the boundary reflection algebra in the enlarged Hilbert space is richer.
@@ 2.2.1 Periodic boundary condition
$w_{2n}=\kappa,\qquad w_{2n+1}=-\kappa.$ (22)
The transfer matrix (5) is
$t(u)=\text{tr}_{0}\Big{(}R_{0L}(u-\kappa)R_{0,L-1}(u+\kappa)\dots R_{02}(u-\kappa)R_{01}(u+\kappa)\Big{)}.$ (23)
@@
$U_{i,j}=\check{R}_{i,j}(2\kappa)=P_{i,j}R_{i,j}(2\kappa).$ (24)
@@
$t(\kappa)=$ $P_{1,L}P_{1,L-1}\dots P_{1,3}P_{1,2}U_{2,3}\dots U_{L-4,L-3}U_{L-2,L-1}U_{L,1}$ (25)
@@
$M=t(-\kappa)^{-1}t(\kappa)=\mathbb{U}^{o}\mathbb{U}^{e},$ (27)
@@
$\mathbb{U}^{o}=\prod_{k=1}^{L/2}U_{2k-1,2k},$ $\mathbb{U}^{e}=\prod_{k=1}^{L/2}U_{2k,2k+1}.$ (28)
The operator $M$ describes one period of Floquet dynamics composed of two steps: $\mathbb{U}^{e}$ and $\mathbb{U}^{o}$. $\mathbb{U}^{e}$ updates every pair of consecutive sites with the first gate starting to act from even positions, while $\mathbb{U}^{o}$ with the first gate starting to act from odd positions [^7] We use the convention to enumerate the sites from right to left., see Fig. 3.
@@
This construction builds an integrable quantum circuit. In fact, since $[t(u,\mathbf{w}),t(v,\mathbf{w})]=0,$ fixing the inhomogeneities to (22), it is easy to see that
$[M,t(z)]=0,$ (29)
with $t(z)$ given in (23).
@@ 2.2.2 Open boundary conditions
$w_{1}=w_{3}=w_{5}=\dots=w_{L}=\kappa,$ $w_{2}=w_{4}=\dots=w_{L-1}=-\kappa.$ (30)
The transfer matrix (16) is now
$t(u)=\text{tr}_{0}\Big{(}{K}_{0}^{L}(u)R_{0L}(u\!-\!\kappa)\dots R_{02}(u\!+\!\kappa)R_{01}(u\!-\!\kappa)K_{0}^{R}(u)R_{10}(u\!+\!\kappa)R_{20}(u\!-\!\kappa)\dots R_{L0}(u\!+\!\kappa)\Big{)}.$ (31)
@@
$t(\kappa)=K_{1}^{R}(\kappa)U_{23}U_{45}\dots U_{L-1,L}U_{12}U_{34}\dots U_{L-2,L-1}\bar{K}_{L}^{L}(\kappa),$ (32)
@@
$\bar{K}_{L}^{L}(\kappa)=\text{tr}_{0}\big{(}K_{0}^{L}(\kappa)R_{0L}(2\kappa)P_{0L}\big{)}.$ (33)
@@
$M=t(\kappa)=\mathbb{U}^{e}\mathbb{U}^{o},$ (34)
@@
$\mathbb{U}^{o}=\bar{K}_{L}^{L}(\kappa)\left(\prod_{k=1}^{\frac{L-1}{2}}U_{2k-1,2k}\right)\,,$ $\mathbb{U}^{e}=\left(\prod_{k=1}^{\frac{L-1}{2}}U_{2k,2k+1}\right)K^{R}_{1}(\kappa).$ (35)
@@ 3 Our construction
In this section, we first discuss in 3.1, the construction of the quantum circuit we are going to study and we clarify the main question we would like to address. In 3.2, we discuss in details the steps that we performed.
@@ 3.1 Building the quantum circuit
$U_{ik,jl}=\check{R}_{ik,jl}(u)=\check{r}_{ij}(u)\check{r}_{kl}(-u).$ (37)
We represent this "double gate" as in Fig. 5, [38].
@@
For open boundary condition, a natural generalization of the right and left reflection algebra in the enlarged space is to consider $k^{R/L}(u)$ and $\tilde{k}^{R/L}(u)$ as two solutions of the boundary Yang-Baxter equation (10) and (11) for $r(u)$ and construct the solution of the boundary Yang-Baxter equation for $R(u)$ as $K^{R/L}_{ij}(u)=k^{R/L}_{i}(u)\tilde{k}^{R/L}_{j}(-u)$, see Appendix A for a proof of this statement. The two-steps Floquet circuit for this construction is Figure 7.
@@ 3.2 Key steps
  • Solve the reflection equation (10) for $K^{R}(u)$;
  • Solve the reflection equation [^8] This solution can be found by applying to the $K^{R}$ solution the automorphism (21). If the $r(u)$ of the model analyzed obeys the crossing unitarity property, this automorphism reduce to the three (equivalent) given in section 2.1.2. (11) for $K^{L}(u)$;
@@
To solve the equation (10), we consider the most general form for the $K^{R}$-matrix, as a $4\times 4$ matrix with entries being functions dependent on a spectral parameter $u$. We label the entries of the $K^{R}$ matrix $\mathcal{K}_{i,j}(u)$, with $i,j$ identifying the row or the column where this element is placed. The expression of the $R$-matrix depends on the model we are analyzing. We normalize [^9] Two $K$ matrices that differ only by a normalization factor solve the same boundary YBE. We choose $\mathcal{K}_{2,2}=1$ since we require that $K^{R}(0)=\mathbb{I}$. Choosing to normalize an off-diagonal element to one won’t be compatible with the regularity condition. the element $\mathcal{K}_{2,2}$ to one. A possible strategy to solve the boundary Yang-Baxter equation is to use a differential approach called Abel’s method. For a review of the method, see [39]. This method consists of taking the formal derivatives of the equation (10) with respect to the spectral parameters $u$, then evaluate it at zero and use the regularity condition
$K^{R}(0)=\mathbb{I},$ $\mathcal{K}_{i,j}(0)=\delta_{i,j}.$ (38)
@@
We take the formal derivative of the boundary Yang-Baxter equation (10),
@@
In this way, we obtain a system of equations depending on the variables $\mathcal{K}_{i,j}(v)$ and $\kappa_{i,j}$. The advantage to solve this system, compared to the initial one given by (10), is that now the equations are linear in the variables. To find all the solutions, we treat $\mathcal{K}_{i,j}{(v)}$ as independent from $\kappa_{i,j}$. We solved some of the equations [^10] The choice on how many equations to solve in the first place depends on the model considered. So far, we did not find a systematic way to make this choice. for some of the variables and plugging back into the remaining equations we solved for the others. Last, we fixed the remaining unknown functions by imposing the compatibility conditions that ${\mathcal{K}_{i,j}}^{\prime}(0)=\kappa_{i,j}$. After having found all the solutions, we discard the ones leading to incompatible conditions [^11] An example of incompatible conditions may be that the element $\mathcal{K}_{2,2}(v)=3$ and $\kappa_{2,2}=5$.. We also discard the solutions that are not compatible with the regularity conditions.
This method is general and only requires that the functions $\mathcal{K}_{i,j}(v)$ are differentiable. This condition, for our case, is not restrictive at all since, as explained in section 5, the $K^{R}$ is one of the building block of the transfer matrix and hence, its derivatives correspond to the conserved charge characterizing the integrable model. Furthermore, (10) contains both the $u$ and $v$, spectral parameter dependence, taking the derivative w.r.t. $u$ and later sending it to zero, won’t imply that we are finding solutions only at order linear in $u$.
This method was vastly used in the literature, in particular Lima-Santos and collaborators, classified the $K$-matrices associated with non-exceptional Lie algebras and superalgebras, see for example [40, 41].
@@ 4.1.1 Heisenberg (XXX) spin chain
We solve the boundary Yang-Baxter equation (10) for $K^{R}$(u) and
$R_{ij,kl}(u)=\big{(}{p}_{ik}+u\mathbb{I}\big{)}\big{(}{p}_{jl}-u\mathbb{I}\big{)}.$ (41)
We solved the equation (10), as explained in the previous section, by using the Mathematica software and we obtained all the solutions for the right reflection matrix $K^{R}(u)$. After detailed investigation, we found that all the solutions of the boundary YBE for two copies of the Heisenberg spin chain take the factorized form
${K}_{ij}^{R}(u)=k^{R}_{i}(u)\tilde{k}^{R}_{j}(-u).$ (42)
Both $k_{i}(u)$ and $\tilde{k}_{j}(u)$ are solutions [^13] We identify as $\tilde{k}$ the $K$-matrix acting on the site $j$ since this solution can be independent from $k(u)$. See Appendix A for details. of the boundary YBE (10) for $r(u)$.
The Heisenberg spin chain is widely studied, diagonal solutions were classified in [3] and general solutions in [42]. For completeness, we write here [^14] This solution is related to the one of [42] by a reparametrization. the most general solution of the reflection algebra for the XXX spin chain with $k^{R}$ regular,
$k^{R}(u)=\left(\begin{array}[]{cc}\frac{\kappa_{1}u+1}{1-\kappa_{1}u}&\frac{\kappa_{2}u}{1-\kappa_{1}u}\\ \frac{\kappa_{4}u}{1-\kappa_{1}u}&1\\ \end{array}\right),$ (45)
where $\kappa_{1}$, $\kappa_{2}$ and $\kappa_{4}$ are free constants. To construct the most general $K^{R}(u)$ from (42) , one can assume that $\tilde{k}_{j}(u)$ has the same form as (45) and relabel the constants $\kappa_{i}$ as $\tilde{\kappa}_{i}$ since now they can in principle take different values. This is described in more details in the Appendix A.
To obtain the solution of the left boundary $K^{L}$ one can use the automorphism (21). This guarantees that if the solution for $K^{R}$ can be written in the factorized form, then the one for $K^{L}$ also factorizes. The $R$-matrix $R(u)$ (41) lacks the crossing unitarity property, thus the solution for $K^{L}$ can be obtained by employing the automorphism (21). The reason is that, for the Heisenberg spin chain, the crossing parameter for $r(u)$ is $\eta=1$, so the crossing unitarity property is broken for $R(u)=r(u)\otimes r(-u)$. However, since we know from the more general automorphism that the solutions for $K^{L}$ should also factorize, we can separately obtain the solution for $k^{L}(u)$ and $\tilde{k}^{L}(u)$. In this way, we obtain
$K^{L}(u)=(k^{R})^{t}(-u-1)\otimes(\tilde{k}^{R})^{t}(u-1).$ (46)
@@ 4.1.2 XXZ spin chain
$r(u)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\sin(u)\csc(u-i\gamma)&-i\sinh(\gamma)\csc(u-i\gamma)&0\\ 0&-i\sinh(\gamma)\csc(u-i\gamma)&\sin(u)\csc(u-i\gamma)&0\\ 0&0&0&1\\ \end{array}\right),$ (51)
@@
$k(u)=\left(\begin{array}[]{cc}\frac{\kappa_{1,1}\sin u+2\cos u}{2\cos u-\kappa_{1,1}\sin u}&\frac{2\kappa_{1,2}}{2\csc u-\kappa_{1,1}\sec u}\\ \frac{2\kappa_{2,1}}{2\csc u-\kappa_{1,1}\sec u}&1\\ \end{array}\right)$ (55)
and $\tilde{k}(u)$ is the same by substituting $\kappa$ with $\tilde{\kappa}$. This solution reproduces (45), in the same spirit as one gets the rational $R$-matrix of the Heisenberg spin chain from the trigonometric solution for the XXZ spin chain, [33]. As observed in [42], the solution of the right boundary Yang Baxter equation is independent from the parameter $\gamma$ (or equivalently $\Delta$). However, from the automorphisms (21) it is clear that $K^{R}$ is connected to $K^{L}$ via the $R$-matrix and hence the anisotropy $\Delta$ (or $\gamma$) of the spin chain enters. We can repeat the same argument of the Heisenberg spin chain, and we obtain
$K^{L}(u)=(k^{R})^{t}(-u+i\gamma)\otimes(\tilde{k}^{R})^{t}(u+i\gamma).$ (56)